OpenStudy (anonymous):

The length of a rectangle is twice its width. If the area of the rectangle is 72 , find its perimeter.

OpenStudy (anonymous):

To get started, relate the length and width. We know that the length is twice the width so l = 2w We know the area is 72, so we use the area formula for a rectangle: lw = 2w(w) = 2w2 =72 Solving for w: 2w2 = 72 w2 = 36 w = 6 The length equals l = 2w l = 12 and the dimensions are 6 x 12 The perimeter is the sum of the lengths of all four sides or 2w + 2l = 2(6) + 2(12) = 36

OpenStudy (shane_b):

\[A=LW=72\]You need to find L and W to get the perimeter since \[P=2L+2W\]To find L and W, rewrite the area equation in terms of one variable knowing that L=2W:\[A=(2W)W\]You should see that W= 6. Now you can use the area equation to solve for L. Once you have L and W, you can find the perimeter quite easily.